[palatino] [12.0pt]             \mr: Ag

\tf \sc \sl \it \bf \bs \bi \tfx \tfxx \tfa \tfb \tfc \tfd

\rm Ag Ag Ag Ag Ag Ag Ag Ag Ag Ag Ag Ag Ag
\ss Ag Ag Ag Ag Ag Ag Ag Ag Ag Ag Ag Ag Ag

\tt Ag  Ag  Ag  Ag  Ag  Ag  Ag  Ag   Ag   Ag  Ag Ag Ag